OpenStudy (kayla2001):

find the slope of a line that is perpendicular to the lie containing the points (-2,-1) and (2,-3).

OpenStudy (anonymous):

do you know the equation to find the slope?

OpenStudy (kayla2001):

do I multiply -3 time -1?

OpenStudy (anonymous):

m= y2-y1 ----- x2-x1 that is the formula to find the slope

OpenStudy (kayla2001):

I don't really understand the process of working out this problem

OpenStudy (anonymous):

okay hold on, i will explain

OpenStudy (anonymous):

first you need to find the slope of the the two given points (-2,-1) (2,-3)

OpenStudy (anonymous):

Do so by using the formula m= y2-y1 / x2-x1

OpenStudy (kayla2001):

@bibby I don't understand

OpenStudy (kayla2001):

my answer was negative 4

OpenStudy (bibby):

find the slope, recall that perpendicular lines have slopes that are negative reciprocals of the original slopes just follow what Quxxn is saying

OpenStudy (kayla2001):

but wont it be -1

OpenStudy (anonymous):

-2/4 sorry i wrote the wrong answer

OpenStudy (kayla2001):

so my anwer would be -1/2

OpenStudy (bibby):

\((-2,-1) and (2,-3).\huge\longrightarrow\frac{-3--1}{2--2}=\frac{-3+1}{2+2}=\frac{-2}{4}\)

OpenStudy (anonymous):

yes, that is the slope of your first equation

OpenStudy (bibby):

the original m is now -1/2 what would the perpendicular lines slope be?

OpenStudy (kayla2001):

-1?

OpenStudy (bibby):

"recall that perpendicular lines have slopes that are negative reciprocals of the original slopes"

OpenStudy (kayla2001):

why did you change your picture I thought you were a girl

OpenStudy (bibby):

rofl, I'm anything you want me to be

OpenStudy (kayla2001):

wait so whats the anwer then we can chit chat

OpenStudy (bibby):

either way, reciprocals are defined as \(\large a \longrightarrow \frac{1}{a}\) so the negative reciprocal of (let's take a random number) 3 is \(\large -\frac{1}{3}\)

OpenStudy (bibby):

find the negative reciprocal of the original lines slope